### v2

@@ 6.2 Issue of LLGRBs
This leads us to another pertinent question about the true nature and origin of LLGRBs. Research by Coward (2005); Liang et al. (2007a); Bernardini et al. (2007); Virgili et al. (2009); Patel et al. (2023) suggests that LLGRBs represent a distinct class of phenomena. Due to their low luminosities, these events are typically observed at low redshifts, such as GRB 060218 at $z=0.033$ (Mirabal et al., 2006) and GRB 980425 at $z=0.0085$ (Tinney et al., 1998). Coward (2005) simulated the local density rate of LLGRBs to be 220 Gpc-3yr-1, based on the under-luminous GRB 980425. Under the assumption that the luminosity of GRB 980425 was typical of such a population of LLGRBs, the authors claim the luminosity cut-off between LLGRBs and high-luminosity GRBs (HLGRBs) to be $2\times 10^{-53}$erg s-1. However, based on GRB 980425 and GRB 060218 Liang et al. (2007a) defined LLGRBs as GRBs with luminosity $<$ 10-49erg s-1. The authors found the LLGRB event rate to be $\sim 325^{+352}_{-177}\,\rm Gpc^{-3}yr^{-1}$ while that of HLGRBs to be $1.12^{+0.43}_{-0.2}$ Gpc-3 yr-1. They also showed that the luminosity functions of LLGRBs and HLGRBs are different and cannot be represented by a single function. Chapman et al. (2007) also investigated the local density rate of LLGRBs based on GRB 980425 and GRB 060218 and found a result of $700\pm 360$ Gpc-3 yr-1 within a distance of 155 Mpc. Virgili et al. (2009) provided further evidence that LLGRBs are a distinct class of GRBs from HLGRBs. They demonstrated that a simple power-law LF is insufficient and that a smooth broken power-law fit is necessary to obtain a good prediction for the distribution of a combined sample of LLGRBs and HLGRBs. Using MC simulation and the LF, they derive an event rate of $\sim 100-400$ Gpc-3 yr-1 for LLGRBs and $\sim 1$ Gpc-3 yr-1 for HLGRBs. However, they conclude that the sample is too small and more data is needed to draw stronger conclusions. Dainotti et al. (2024b) obtained a GRB event rate of $8.47-9$ Gpc-3 $yr^{-1}$ between $2.31.9<z<$ 2.3 using $z_{obs}$. This matches the HLGRB rate found by Dong et al. (2023); Liang et al. (2007a); Virgili et al. (2009). Lan et al. (2021) utilized 5 LLGRBs out of a collection of $1$ LGRBs obtained from Swift to demonstrate that a triple power-law LF provides a superior fit for the data, again indicating the existence of LLGRBs as a separate component. Recently, Dong et al. (2023) demonstrated that the event rate of HLGRB closely follows the SFR, while that of LLGRB shows an excess at low-$z$ ($z<1$). Combining the sample leads to the observed low-$z$ excess and mismatch between the GRBFR and SFR. These previous works strongly suggest that LLGRBs are a distinct class of GRBs with the potential to solve the GRB-SFR discrepancy. However, to draw more definitive conclusions, a larger sample of GRBs with redshift is needed.
6.3 Discrepancy with SFR
@@ 7 Summary, and Conclusion
In this study, we have built upon the methodology established by Dainotti et al. (2024b) and Dainotti et al. (2024c). We use a 20% larger training set and more robust formula estimation to predict the redshifts of 276 GRBs. In addition, we perform an extensive MC simulation to evaluate how our ML model will perform in the upcoming years. For the benefit of the community, we have a publicly available and free-to-use web app designed to provide pseudo-redshifts for any GRB with plateau emission, along with additional functionalities.
We draw the following conclusions from this research:
@@
  • The ML results obtained in this analysis, presented in Sec. 5.1, have similar accuracy with those presented in Dainotti et al. (2024c) but slightly lower. Specifically, we have a 10% decrease in $r$, from 0.719 to 0.644, and a 9.8% increase in RMSE, from 0.91 to 1. These changes in the metrics can be attributed to two reasons. First, the new sample of GRBs added to the training set has a higher percentage of GRBs at larger redshift, as compared to the training set in Dainotti et al. (2024c) (see Sec. LABEL:comparesets). This can lead to the ML model having to adjust to the new data with a larger intrinsic scatter, thus leading to a decrease in the performance metrics. The second reason is the change in the methodology for formula estimation. In Dainotti2024GRBRedshift, the formulas estimated for GAM and GLM were determined using a single training-test split instance. Here, we have expanded the methodology and generalized the formulas. Generalizing the formula here means it can better capture the underlying relations among GRB properties for a larger population of GRBs rather than just a small, limited sample of the training set. However, this generalization leads to poorer performance on a singular training-test split as the variability of the underlying relation is higher in a smaller sample. However, the formulas generalize well to additional synthetic GRB samples with a more uniform redshift distribution, which is evident from the improved balanced sampling result across all the metrics (lower right and left panels of Fig. 13). This promising result indicates that our ML model has effectively generalized to the entire GRB sample and the MC simulations further emphasize this point.
@@
  • In Section 5.2, we demonstrated that $\log(NH)$ exerts the greatest influence on redshift predictions. This finding is consistent with the results reported in Dainotti et al. (2024c) and Dainotti et al. (2024b). A similar result was also discovered independently by Racz (2023), where they found that the performance of their ML model was reduced when they removed $\log(NH)$ from their list of predictors. We also find that $\log(T_{a})$ is the second most important predictor of redshift, similar to Dainotti et al. (2024c), whereas Dainotti et al. (2024b) found that $\log(F_{a})$ is the second most predictive feature. Given the correlation of $\log(T_{a})$ and $\log(F_{a})$, these results demonstrate that plateau features are important when predicting the redshift of GRBs.
@@
  • In Sec 5.3, we predict the redshift of 276 GRBs using our trained model. Both the bias-corrected and non-bias-corrected predictions are freely available to the community through this publication. However, we still have a discrepancy between the predicted redshift distributions and the redshift distribution of the training set. Dainotti et al. (2024c) questioned as to what extent the variation between the redshift distributions of GRBs with and without $z_{obs}$ are inherently different, or if the variation might be attributed to the limited size of the training set. In this analysis, we have increased the training set with 30 additional GRBs. However, we still need to investigate the origin of the different distributions. A KS-test comparing the $\log(NH)$ distributions for the training and the generalization sets still presents us with the conclusion that they are derived from different parent populations. Given the fact that $\log(NH)$ is the most important predictor, this could explain the discrepancy in the redshift distributions.
@@
  • The MC simulations of future data in Sec. 5.4 present us with promising results. Based on those estimates, we conclude that our ML model is robust enough to last for at least three more years before major retraining is needed. The histogram distributions for $r$, RMSE, and MAD show that our ML model performs better when more sources are added to the training set. It’s important to note that this condition holds only if the additional data incorporated into the training set in the future comes from the same parent population as the existing sample.
@@
  • The methodology described here can be accessed via the associated web app. It enables users to obtain redshift predictions based on the ML model presented here for their GRBs. It also enables users to choose from a set of predefined models that, in our experience, work well with GRB samples and train these models with their data. Then, they can obtain redshift predictions from this custom model. The web app is designed to provide results like the ones presented in this work for any custom data the user may input.
@@
  • In Sec 5.5, we presented calculations on how many GRB’s pseudo-redshifts we will be able to obtain in the future using SVOM data, and possibly THESEUS. The numbers mentioned there are based on current estimates from publications of SVOM and Swift’s observational statistics. These numbers will need to be updated as more GRBs by SVOM are detected. Regardless, the numbers paint a promising picture of the future applicability of such ML techniques. In fact, this work has already enabled us to increase the Swift GRB sample with redshift by $\sim$110%.
8 Acknowledgments
@@
The authors thank Omar Elsherif for his contributions to building the basic structure of the webapp, upon which the whole functionality is based. We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2024/017205. We acknowledge Biagio de Simone and Christopher Cook for helping with the MC computations. We acknowledge funding from Swift GI Cycle XIX, grant number 22-SWIFT22-0032.
9 Appendix
@@
Here, we are presenting the histogram distributions of the training set from Dainotti et al. (2024c) and the training set used in this work. The Fig. 19 shows the previous sample with solid lines and the new sample with dashed lines. Each feature has the KS-test and AD-test statistics mentioned at the top of the histogram. As mentioned in Sec 3, $T_{90}$, and $\gamma$ obtain p-values $<$ 0.05 in both AD-test and KS-test, while $\log(F_{a})$ obtains p-value $<$ 0.05 only in AD-test. The case can be explained by looking at the distribution, where we see a heavy-tailed distribution for the newer data set (dashed lines). Since the AD-test is more sensitive to heavy-tailed distributions, we obtain a p-value of $<$ 0.05.
@@
Figure 19: The histogram distribution comparing the current training set with the previous training set from Dainotti et al. (2024c). At the top of each distribution, the KS-test and AD-test comparison results are shown.
@@
The weights assigned by SuperLearner to its constituent models vary depending on the data set it is being used. Fig.20 shows the weight assigned to EARTH is lower than those of Random forest, GAM, and GLM models. However, in Narendra et al. (2022), EARTH was the highest-weighted model. Thus, we can say that our final ensemble model is well-suited for the size of our data sample.
Figure 20: Barplot showing the weights assigned by SuperLearner to an ensemble of GAM, GLM, Random forest and EARTH models.

### v3

@@ 6.2 Issue of LLGRBs
This leads us to another pertinent question about the true nature and origin of LLGRBs. Research by Coward (2005); Liang et al. (2007a); Bernardini et al. (2007); Virgili et al. (2009); Patel et al. (2023) suggests that LLGRBs represent a distinct class of phenomena. Due to their low luminosities, these events are typically observed at low redshifts, such as GRB 060218 at $z=0.033$ (Mirabal et al., 2006) and GRB 980425 at $z=0.0085$ (Tinney et al., 1998). Coward (2005) simulated the local density rate of LLGRBs to be 220 Gpc-3yr-1, based on the under-luminous GRB 980425. Under the assumption that the luminosity of GRB 980425 was typical of such a population of LLGRBs, the authors claim the luminosity cutoff between LLGRBs and high-luminosity GRBs (HLGRBs) to be $2\times 10^{-53}$erg s-1. However, based on GRB 980425 and GRB 060218 Liang et al. (2007a) defined LLGRBs as GRBs with luminosity $<$ 10-49erg s-1. The authors found the LLGRB event rate to be $\sim 325^{+352}_{-177}\,\rm Gpc^{-3}yr^{-1}$ while that of HLGRBs to be $1.12^{+0.43}_{-0.2}$ Gpc-3 yr-1. They also showed that the LFs of LLGRBs and HLGRBs are different and cannot be represented by a single function. Chapman et al. (2007) also investigated the local density rate of LLGRBs based on GRB 980425 and GRB 060218 and found a result of $700\pm 360$ Gpc-3 yr-1 within a distance of 155 Mpc. Virgili et al. (2009) provided further evidence that LLGRBs are a distinct class of GRBs from HLGRBs. They demonstrated that a simple power-law LF is insufficient and that a smooth broken power-law fit is necessary to obtain a good prediction for the distribution of a combined sample of LLGRBs and HLGRBs. Using MC simulation and the LF, they derive an event rate of $\sim 100-400$ Gpc-3 yr-1 for LLGRBs and $\sim 1$ Gpc-3 yr-1 for HLGRBs. However, they conclude that the sample is too small and more data are needed to draw stronger conclusions. Dainotti et al. (2024b) obtained a GRB event rate of $8.47-9$ Gpc-3 $yr^{-1}$ between $2.31.9<z<$ 2.3 using $z_{obs}$. This matches the HLGRB rate found by Dong et al. (2023); Liang et al. (2007a); Virgili et al. (2009). Lan et al. (2021) utilized 5 LLGRBs out of a collection of $1$ LGRBs obtained from Swift to demonstrate that a triple power-law LF provides a superior fit for the data, again indicating the existence of LLGRBs as a separate component. Recently, Dong et al. (2023) demonstrated that the event rate of HLGRB closely follows the SFR, while that of LLGRB shows an excess at low-$z$ ($z<1$). Combining the sample leads to the observed low-$z$ excess and mismatch between the GRBFR and SFR. These previous works strongly suggest that LLGRBs are a distinct class of GRBs with the potential to solve the GRB-SFR discrepancy. However, to draw more definitive conclusions, a larger sample of GRBs with redshift is needed.
6.3 Discrepancy with star formation rate
@@ 7 Summary and conclusion
In this study, we have built upon the methodology established by Dainotti et al. (2024b) and Dainotti et al. (2024c). We used a 20% larger training set and more robust formula estimation to predict the redshifts of 276 GRBs. In addition, we performed an extensive MC simulation to evaluate how our ML model will perform in the upcoming years. For the benefit of the community, we have produced a publicly available and free-to-use web app designed to provide pseudo-redshifts for any GRB with plateau emission, along with other additional functionalities. From this reasearch, we draw the following conclusions:
@@
  • The ML results obtained in this analysis, presented in Sec. 5.1, have a similar accuracy with those presented in Dainotti et al. (2024c) but are slightly lower. Specifically, we have a 10% decrease in $r$, from 0.719 to 0.644, and a 9.8% increase in RMSE, from 0.91 to 1. These changes in the metrics can be attributed to two reasons. First, the new sample of GRBs added to the training set has a higher percentage of GRBs at a larger redshift, as compared to the training set in Dainotti et al. (2024c) (see Sec. 3). This can lead to the ML model having to adjust to the new data with a larger intrinsic scatter, thus leading to a decrease in the performance metrics. The second reason is the change in the methodology for formula estimation. In Dainotti et al. (2024c), the formulas estimated for GAM and GLM were determined using a single training-test split instance. Here, we expanded the methodology and generalized the formulas. Generalizing the formula here means it can better capture the underlying relations among GRB properties for a larger population of GRBs rather than just a small, limited sample of the training set. However, this generalization leads to poorer performance on a singular training-test split, as the variability of the underlying relation is higher in a smaller sample. Nonetheless, the formulas generalize well to additional synthetic GRB samples with a more uniform redshift distribution, which is evident from the improved balanced sampling result across all the metrics (lower-right and -left panels of Fig. 13). This promising result indicates that our ML model has effectively generalized to the entire GRB sample, and the MC simulations further emphasize this point.
@@
  • In Section 5.2, we demonstrated that $\log(NH)$ exerts the greatest influence on redshift predictions. This finding is consistent with the results reported in Dainotti et al. (2024c) and Dainotti et al. (2024b). A similar result was also discovered independently by Racz (2023), who found that the performance of their ML model was reduced when they removed $\log(NH)$ from their list of predictors. We also find that $\log(T_{a})$ is the second most important predictor of redshift, similar to Dainotti et al. (2024c). However, we note that Dainotti et al. (2024b) found that $\log(F_{a})$ is the second most predictive feature. Given the correlation of $\log(T_{a})$ and $\log(F_{a})$, these results demonstrate that plateau features are important when predicting the redshift of GRBs.
@@
  • In Sec 5.3, we predicted the redshift of 276 GRBs using our trained model. Both the bias-corrected and non-bias-corrected predictions are freely available to the community through this publication. However, we still have a discrepancy between the predicted redshift distributions and the redshift distribution of the training set. Dainotti et al. (2024c) questioned the extent the variation between the redshift distributions of GRBs with and without $z_{obs}$ are inherently different and whether the variation might be attributed to the limited size of the training set. In this analysis, we have increased the training set with 30 additional GRBs. However, we still need to investigate the origin of the different distributions. A KS test comparing the $\log(NH)$ distributions for the training and the generalization sets still presents us with the conclusion that they are derived from different parent populations. Given the fact that $\log(NH)$ is the most important predictor, this could explain the discrepancy in the redshift distributions.
@@
  • The MC simulations of future data in Sec. 5.4 present us with promising results. Based on those estimates, we conclude that our ML model is robust enough to last for at least three more years before major retraining is needed. The histogram distributions for $r$, RMSE, and MAD show that our ML model performs better when more sources are added to the training set. It is important to note that this condition holds only if the additional data incorporated into the training set in the future comes from the same parent population as the existing sample.
@@
  • The methodology described here can be accessed via the associated web app. The app enables users to obtain redshift predictions based on the ML model presented here for their GRBs. It also enables users to choose from a set of predefined models that, in our experience, work well with GRB samples and to subsequently train the models with their data. Then, they can obtain redshift predictions from the customized model built using the web app. The web app is designed to provide results similar to those presented in this work for any custom data the user may input.
@@
  • In Sec 5.6, we presented calculations on how many GRB’s pseudo-redshifts we will be able to obtain in the future using SVOM data and possibly THESEUS. The numbers mentioned there are based on current estimates from publications of SVOM and Swift’s observational statistics. These numbers will need to be updated as more GRBs by SVOM are detected. Regardless, the numbers paint a promising picture of the future applicability of such ML techniques. In fact, this work has already enabled us to increase the Swift GRB sample with redshift by roughly 110%.
In conclusion, we envision our web app and ML based redshift estimates will help the community in follow-up new GRBs in the coming years and the open source nature of the web app will help foster continuous development and allow new discoveries.
8 Data availability
The web app is available for free on the following Github page:https://github.com/gammarayapp/GRB-Web-App.
Acknowledgements.
@@
The authors thank Omar Elsherif for his contributions to building the basic structure of the web app, upon which the whole functionality is based. We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2024/017205. We acknowledge Biagio de Simone and Christopher Cook for helping with the MC computations. We acknowledge funding from Swift GI Cycle XIX, grant number 22-SWIFT22-0032. Numerical computations were in part carried out on Small Parallel Computers at the Center for Computational Astrophysics, National Astronomical Observatory of Japan.
Appendix A Distribution of training sets
@@
Here, we are presenting the histogram distributions of the training set from Dainotti et al. (2024c) and the training set used in this work. The Fig. 19 shows the previous sample with solid lines and the new sample with dashed lines. Each feature has the KS test and AD test statistics mentioned at the top of the histogram. As mentioned in Sec 3, $T_{90}$, and $\gamma$ obtain p-values $<$ 0.05 in both AD test and KS test, while $\log(F_{a})$ obtains p-value $<$ 0.05 only in AD test. The case can be explained by looking at the distribution, where we see a heavy-tailed distribution for the newer data set (dashed lines). Since the AD test is more sensitive to heavy-tailed distributions, we obtain a p-value of $<$ 0.05.
@@
Figure 19: Distribution comparing the current training set with the previous training set from Dainotti et al. (2024c). At the top of each distribution, the KS test and AD test comparison results are shown.
Appendix B Weight distribution by SuperLearner
@@
The weights assigned by SuperLearner to its constituent models vary depending on the data set it is being used. Fig.20 shows the weight assigned to EARTH is lower than those of RF, GAM, and GLM models. However, in Narendra et al. (2022), EARTH was the highest-weighted model. Thus, we can say that our final ensemble model is well-suited for the size of our data sample.
Figure 20: Weights assigned by SuperLearner to an ensemble of GAM, GLM, RF, and EARTH models.